Find the value of x so the line that passes through (x, 5) and (12, 2) is perpendicular to the line that passes through (0, 1) and (2, 7).
Answer by rwm(914)

slope of second line
(7-1)/(2-0)
slope of first line
(5-2)/(x-12)
we want the two slopes multiplied to equal -1
(7-1)/(2-0)*(5-2)/(x-12)=-1
it looks harder than it is.
just do the math.
6/2*3/(x-12)=-1
9/x-12=-1
9=-1*(x-12)
-9=x-12
add 12
3=x